find the vertex and x-intercept(s) of the function given below y=x^2+4x+7
Answer by ewatrrr(24785)

y=x^2+4x+7
y = (x+2)^2 -4 +7
y = (x+2)^2 + 3, V(-2,3)
0 = (x+2)^2 + 3
x = -2 ± i√3 NO x-intercepts